39. Alberto Trucking is expected to pay an annual dividend of $2.50 per share next year....

39.

Alberto Trucking is expected to pay an annual dividend of $2.50 per share next year. The stock is currently selling for $33.50 a share. What is the expected total return on this stock if that return is equally divided between a dividend yield and a capital gains yield?

11.19 percent 14.93 percent 3.73 percent 7.46 percent

40.

Rates of Return An investor earned an arithmetic average return over 4 years of 14.3%. The first year's return was 14.8%, the second year's return was -1.40%, the third year's return was 27.4%. What was the fourth year's return?

18.36%

16.40%

14.30%

19.17%
Answer #1

1.Dividend yield=Dividend/Current price

(2.5/33.5)=7.46%(Approx).

Dividend yield=Capital gains yield=7.46%(Approx)

Total return=Dividend yield+Capital gains yield

=(7.46+7.46)=14.93%(Approx).

2.

arithmetic average return=Total return/Total time period

14.3=(14.8-1.4+27.4+fourth year's return)/4

(14.3*4)=40.8+fourth year's return

fourth year's return=57.2-40.8

=16.4%